# RoboBrain: A Unified Brain Model for Robotic Manipulation from Abstract to Concrete – Supplementary Material –

### v1

@@ Appendix A Details of Models and Training
Model Setting. RoboBrain is built upon the LLaVA [LLaVa] framework and consists of three main components: the visual encoder, projector, and large language model (LLM).
For the visual encoder, we utilized the SigLIP [siglip] model, specifically the siglip-so400m-patch14-384, which is pre-trained on WebLi [pali] at a resolution of 384x384. The SigLIP model improves upon traditional CLIP [clip] architectures by employing a sigmoid loss function that operates solely on image-text pairs, eliminating the need for global normalization of pairwise similarities. This enhancement allows for more efficient scaling of batch sizes while maintaining performance, even at smaller scales. SigLIP has 27 hidden layers and processes input images using patches of size 14x14, resulting in 729 visual tokens per image. The projector consists of a two-layer MLP [llavanext] that projects the visual tokens obtained from the visual encoder to the dimensions of the text embeddings. For the LLM, we adopted the Qwen2.5-7B-Instruct [qwen2_5] model, which is a state-of-the-art open-source LLM that is part of the latest Qwen series [Qwen]. It features 28 hidden layers and supports long-context inputs of up to 128K tokens, providing multilingual capabilities across more than 29 languages.
In Stage 4, we introduced LoRA [lora] to train RoboBrain, enabling it to acquire affordance perception and trajectory prediction capabilities. LoRA is a technique that allows for parameter-efficient fine-tuning of large models by adding low-rank parameter matrices to existing layers. We incorporated LoRA modules with a rank of 64 into the feed-forward network layers of both the Projector and the LLM, freezing all parameters except those of the LoRA modules during training.
Training Setting. In the main text of the paper, we employed a staged training strategy, with complete settings presented in Tab. 1. We primarily referenced the training strategy of LLaVA-Onevision [li2024llava], a state-of-the-art multimodal large language model, and built upon this foundation to expand the robotic training phase. During the entire training phase, we conducted all experiments on a cluster of servers, each equipped with 8$\times$×A800 GPUs.
@@
[…]
@@ Appendix B Details of Training Dataset
  • LCS-558K is a subset of the LAION/CC/SBU dataset [cc3m, laion], specifically designed as the LLaVA Visual Instruct Pretrain [LLaVa] Dataset. This dataset has been carefully filtered to achieve a more balanced distribution of concept coverage, ensuring diverse and representative visual content. The primary purpose of LCS-558K is to facilitate the alignment between the visual encoder and the LLM, enabling the LLM to comprehend visual information.
@@
  • Image-4M comprises 8 data sources, including 3 from the LLaVA-Recap series [li2024llava-next]: BLIP558K, COCO118K, and CC3M, as well as UReader [ye2023ureader], Instruct Azure DC [li2024llava-next], Evol-Instruct [chen2024allava], and SynthDog [kim2022donut] We utilized the download links provided by the LLaVA-OneVision team for the data acquisition.
@@
  • SI-3.2M [^1] Due to the unavailability of certain datasets, the actual data used amounts to 3.1M. [li2024llava] consists of 3.2 million samples, carefully curated to support multimodal learning. It includes subsets from existing datasets such as Cambrian [tong2024cambrian], Cauldron [li2024llava-next], and UReader [ye2023ureader], which were subjected to cleaning and re-annotation to ensure data quality. Additionally, it incorporates single-image data from sources like AI2D [ai2d] and OKVQA [marino2019okvqa], alongside a newly compiled single-image collection designed to achieve a balanced and diverse dataset.
@@
  • OV-1.6M [^2] Due to the vague descriptions and missing key information regarding dataset filtering in the original paper, we ended up using 2.4M data. [li2024llava] comprises 1.6 million samples, which includes approximately 800K high-quality samples re-sampled from earlier SI-3.2M datasets with a data replay strategy, ensuring improved data reliability and relevance. Additionally, the dataset incorporates M4-Instruct data to enrich instructional learning tasks. A significant component of OV-1.6M is its video data, which has been released alongside LLaVA-video data. The video subset used in the dataset is specifically aligned with the previous annotation format, providing a diverse multimodal resource for advancing vision-language learning.
@@
  • RoboVQA-800K [RoboVQA] consists of realistic data gathered from various user requests, utilizing different embodiments including robots, humans, and humans equipped with grasping tools. The dataset features 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, with each episode accompanied by corresponding image and text prompt inputs. The primary purpose of RoboVQA-800K is to enhance RoboBrain’s reasoning capabilities in robotic-related scenarios.
@@
  • ScanView-318K totals 318K samples, which integrates data from several high-quality sources, including MMScan-224K [mmscan], 3RScan-43K [3rscan], ScanQA-25K [scanqa], and SQA3D-26K [sqa3d], each contributing unique strengths. MMScan-224K provides multimodal scene data with detailed annotations, such as object segmentation and textual descriptions. 3RScan-43K offers 3D reconstructions and semantic annotation. ScanQA-25K includes question-answer pairs based on 3D scanned environments. SQA3D-26K focuses on spatial question answering. Together, these datasets provide diverse scene-scanning image data, long video sequences, and high-resolution samples, equipping models with fine-grained environmental perception and reasoning abilities.
@@ Appendix C Complementary Experiments
Additionally, we explore the impact of different architectures and pre-trained VLMs, as well as different LLM backbones on our experimental results. We also conduct ablation studies at various stages to meticulously analyze the contributions of each stage to overall performance.
@@
Dataset Split RoboBrain (Ours) GPT-4V [gpt4v] LLaVA-OV-7B [llavanext] InternVL2-8B [internvl] Qwen2-VL-7B [qwen2vl] GPT-4o [gpt4o]
A12D[ai2d] test 82.03 78.2 81.4 83.8 - 94.2
ChartQA[chartqa] test 80.48 78.5 80 83.3 83 85.7
DocVQA[docvqa] test 88 88.4 87.5 91.6 94.5 92.8
TextVQA[textvqa] val 75.85 - 71.07 77.4 84.3 -
MMMU[mmmu] val 49 56.8 48.8 51.8 54.1 69.1
MMStar[mmstar] test 61.23 57.1 61.7 61.5 60.7 63.9
OCRBench[ocrbench] - 677 656 697 794 845 805
RealWorldQA[realworldqa] test 68.89 61.4 66.3 64.4 70.1 58.6
SeedBench[seedbench] image 71.03 49.9 75.4 76.2 - 76.2
MMbench[mmbench] en-dev 81.52 81.3 83.2 - - 83.4
MMbench[mmbench] en-test 80.44 75 80.8 81.7 83 -
MME[mme] test 2084 1926 1998 2210 2327 -
@@ C.1 More Results on General Benchmarks
To evaluate performance on general tasks in real-world scenarios, as is commonly done with MLLMs [gpt4v, llavanext, internvl, qwen2vl, gpt4o], we conducted experiments using a diverse set of image benchmarks summarized in Table 2. We leveraged the comprehensive evaluation toolkit, LMMs-Eval[lmms, zhang2024unveiling], to evaluate RoboBrain’s performance on general benchmarks. These benchmarks are categorized into three classes:
  • Chart, Diagram, and Document Understanding. As key visual formats for structured OCR data, benchmarks such as AI2D [ai2d], ChartQA [chartqa], DocVQA [docvqa], and OCRBench [ocrbench] were utilized. Open-source models like InternVL2-8B [internvl] and LLAVA-OV-7B [llavanext] have demonstrated comparable performance to closed-source models such as GPT-4V [gpt4v]. For RoboBrain, despite being optimized primarily for multidimensional robotic tasks, it surpasses LLAVA-OV-7B [llavanext] and GPT-4V [gpt4v] on these benchmarks, achieving a significant improvement in structured OCR tasks, with the only exceptions being DocVQA [docvqa], where it performs slightly lower than GPT-4V [gpt4v], and OCRBench [ocrbench], where it falls slightly behind LLAVA-OV-7B [llavanext].
  • Visual Perception and Multi-domain Reasoning. This category focuses on complex visual perception and multidisciplinary reasoning tasks. Benchmarks for visual perception include MMStar [mmstar], MMBench [mmbench], and MME [mme], while reasoning benchmarks include MMMU [mmmu] and SeedBench [seedbench]. RoboBrain demonstrates comparable performance to GPT-4V [gpt4v] and LLAVA-OV-7B [llavanext] across multiple benchmarks.
  • Real-world Understanding and Interaction. Evaluating MLLMs [gpt4v, llavanext, internvl, qwen2vl, gpt4o] as general-purpose assistants in real-world settings is crucial, as these scenarios extend beyond controlled environments. For this, the RealworldQA [realworldqa] benchmark was utilized. Results indicate that RoboBrain not only outperforms open-source models like LLAVA-OV-7B [llavanext] and InternVL2-8B [internvl], but also exceeds closed-source models such as GPT-4V [gpt4v] and GPT-4o [gpt4o], showcasing its extensive knowledge base and strong generalization capabilities.
@@
Dataset Split / Metric RoboBrain (Ours) GPT-4V [gpt4v] LLaVA-OV-7B [llavanext] RoboMamba [Robomamba] Qwen2-VL-7B [qwen2vl]
RoboVQA[RoboVQA] BLEU1 72.05 32.23 38.12 54.9 33.22
@@
OpenEQA[OpenEQA] OBJECT-STATE-RECOGNITION 70.4 63.2 72.02 - 72.06
@@
To evaluate RoboBrain’s performance on robotic capabilities in real-world scenarios, we selected RoboVQA [RoboVQA], OpenEQA [OpenEQA], and the test set of ShareRobot, extracted from the proposed ShareRobot dataset, as robotic benchmarks for multi-dimensional assessment, as shown in Table 3. The chosen baselines include MLLMs such as GPT-4V [gpt4v], LLaVA-OV-7B [llavanext], and Qwen2-VL-7B [qwen2vl], as well as robotic models like RoboMamba [Robomamba]. Detailed descriptions of the three selected robotic benchmarks and the analysis of each results are provided below:
@@
  • OpenEQA [OpenEQA] provides a robotics VQA benchmark with over 1,600 high-quality human-generated questions drawn from more than 180 real-world scenes, targeting the task of Embodied Question Answering (EQA) for environment understanding. For fairness, we evaluated all models using the prompt templates and the LLM-Score metric provided by OpenEQA [OpenEQA]. Based on the evaluation results, our proposed model, RoboBrain, outperforms GPT-4V [gpt4v] overall and achieves comparable performance to other baselines. In the future, we plan to further enhance RoboBrain’s spatial intelligence to improve its generalization across scenes.
  • ShareRobot (Eval) provides a cross-scene and cross-embodiment robotics benchmark consisting of 2,050 VQA pairs, drawn from 102 diverse scenes (e.g., bedroom, laboratory, kitchen, office) and covering 12 different robot bodies. Similar to RoboVQA [RoboVQA], we categorized various robotic VQA tasks into planning, planning with context, planning remaining steps, future prediction, generative affordance, past description, success (positive/negative), and discriminative affordance (positive/negative). Unlike RoboVQA benchmark [RoboVQA], we utilized GPT-4o [gpt4o] to score the evaluation results instead of BLEU metrics for each task, aiming for more accurate performance assessment. Based on the results, our proposed model, RoboBrain, outperforms all baselines, demonstrating its exceptional planning capabilities across diverse scenes and embodiments.
@@ C.3 Effectiveness of ShareRobot
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 4. As shown in the table, the inclusion of ShareRobot data enhances the model’s performance compared to scenarios without ShareRobot.
Table 4: Experimental results for effectiveness of ShareRobot and different robot data proportion.
@@
OneVision ShareRobot Others Type-1 Type-2 Type-3 RoboVQA[RoboVQA] OpenEQA[OpenEQA] ShareRobot
@@ C.4 Effectiveness of Robot Data Proportion
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 4. As shown in the table, a 4:6 ratio of robotic data represents a good choice for training data, effectively balancing performance on both the robotic and the general benchmark.
Table 5: Additional Experimental Results. “SFT Data (G:R)” indicates the ratio of training data for fine-tuning VLMs, where “G” represents general VQA data and “R” denotes robot data (with half being ShareRobot). The total dataset size is 1.47M.
[…]
C.5 Different Architecture and Pre-trained VLMs
To validate the effectiveness of different architecture and pre-trained VLMs and training data in the stage 3 training setup, we selected LLaVA-OneVision-7B [li2024llava], OpenVLA-7B [OpenVLA], and Qwen2VL-7B [qwen2vl], each representing a distinct architecture among VLMs, and conducted SFT using the same proportion of training data described in the main text. As shown in Tab. LABEL:tab_additional (a), the results demonstrated that incorporating ShareRobot can significant performance improvements. For unaligned VLMs like LLaVA 1.5 [llava-1.5] and OpenVLA, we first aligned the MLP using BLIP-558k [li2024llava-next]; other models were directly finetuned.

### v2

@@ Appendix A Details of Models and Training
Model Setting. RoboBrain is built upon the LLaVA [48] framework and consists of three main components: the visual encoder, projector, and large language model (LLM).
For the visual encoder, we utilized the SigLIP [92] model, specifically the siglip-so400m-patch14-384, which is pre-trained on WebLi [15] at a resolution of 384x384. The SigLIP model improves upon traditional CLIP [70, 28] architectures by employing a sigmoid loss function that operates solely on image-text pairs, eliminating the need for global normalization of pairwise similarities. This enhancement allows for more efficient scaling of batch sizes while maintaining performance, even at smaller scales. SigLIP has 27 hidden layers and processes input images using patches of size 14x14, resulting in 729 visual tokens per image. The projector consists of a two-layer MLP [44] that projects the visual tokens obtained from the visual encoder to the dimensions of the text embeddings. For the LLM, we adopted the Qwen2.5-7B-Instruct [80] model, which is a state-of-the-art open-source LLM that is part of the latest Qwen series [5]. It features 28 hidden layers and supports long-context inputs of up to 128K tokens, providing multilingual capabilities across more than 29 languages.
In Stage 4, we introduced LoRA [29] to train RoboBrain, enabling it to acquire affordance perception and trajectory prediction capabilities. LoRA allows for parameter-efficient fine-tuning of large models by adding low-rank parameter matrices to existing layers. We incorporated LoRA modules with a rank of 64 into the feed-forward network layers of both the Projector and the LLM, freezing all parameters except those of the LoRA modules during training.
Training Setting. In the main text of the paper, we employed a staged training strategy, with complete settings presented in Tab. 4. We primarily referenced the training strategy of LLaVA-Onevision [41], a state-of-the-art multimodal large language model, and built upon this foundation to expand the robotic training phase. During the entire training phase, we conducted all experiments on a cluster of servers, each equipped with 8$\times$×A800 GPUs.
@@
LR: $\bm{\psi}_{\text{ViT}}$ 2 ×$\times 10^{-6}$ 2 ×$\times 10^{-6}$ 2 ×$\times 10^{-6}$ 2 ×$\times 10^{-6}$ 2 ×$\times 10^{-6}$ 2 ×$\times 10^{-6}$
LR: $\{\bm{\theta}_{\text{Proj.}},\bm{\phi}_{\text{LLM}},\bm{\phi}_{\text{LoRA}}\}$ 1×$\times 10^{-3}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$
@@ Appendix B Details of Training Dataset
  • LCS-558K is a subset of the LAION/CC/SBU dataset [11, 72], specifically designed as the LLaVA Visual Instruct Pretrain [48] Dataset. This dataset has been filtered to ensure a balanced distribution of concept coverage, providing diverse and representative visual content. The primary purpose of LCS-558K is to facilitate alignment between the visual encoder and the LLM, enabling the LLM to understand visual information.
@@
  • Image-4M comprises 8 data sources, including 3 from the LLaVA-Recap series [40]: BLIP558K, COCO118K, and CC3M, as well as UReader [90], Instruct Azure DC [40], Evol-Instruct [13], and SynthDog [36] We utilized the download links provided by the LLaVA-OneVision team for the data acquisition.
@@
  • SI-3.2MDue to the unavailability of certain datasets, the actual data used amounts to 3.1M. [41] consists of 3.2 million samples, carefully curated to support multimodal learning. It includes subsets from existing datasets such as Cambrian [81], Cauldron [40], and UReader [90], which were subjected to cleaning and re-annotation to ensure data quality. Additionally, it incorporates single-image data from sources like AI2D [34] and OKVQA [62], alongside a newly compiled single-image collection designed to achieve a balanced and diverse dataset.
@@
  • OV-1.6MDue to the vague descriptions and missing key information regarding dataset filtering in the original paper, we ended up using 2.4M data. [41] comprises 1.6 million samples, which includes approximately 800K high-quality samples re-sampled from earlier SI-3.2M datasets with a data replay strategy, ensuring improved data reliability and relevance. Additionally, the dataset incorporates M4-Instruct data to enrich instructional learning tasks. A significant component of OV-1.6M is its video data, which has been released alongside LLaVA-video data. The video subset used in the dataset is specifically aligned with the previous annotation format, providing a diverse multimodal resource for advancing vision-language learning.
@@
  • RoboVQA-800K [73] consists of realistic data gathered from various user requests, utilizing different embodiments including robots, humans, and humans equipped with grasping tools. The dataset features 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, with each episode accompanied by corresponding image and text prompt inputs. The primary purpose of RoboVQA-800K is to enhance RoboBrain’s reasoning capabilities in robotic-related scenarios.
@@
  • ScanView-318K totals 318K samples, which integrates data from several high-quality sources, including MMScan-224K [59], 3RScan-43K [83], ScanQA-25K [4], and SQA3D-26K [60], each contributing unique strengths. MMScan-224K provides multimodal scene data with detailed annotations, such as object segmentation and textual descriptions. 3RScan-43K offers 3D reconstructions and semantic annotation. ScanQA-25K includes question-answer pairs based on 3D scanned environments. SQA3D-26K focuses on spatial question answering. Together, these datasets provide diverse scene-scanning image data, long video sequences, and high-resolution samples, equipping models with fine-grained environmental perception and reasoning abilities.
@@ Appendix C Complementary Experiments
Additionally, we explore the impact of different architectures and pre-trained MLLMs, as well as different LLM backbones on our experimental results. We also conduct ablation studies at various stages to meticulously analyze the contributions of each stage to overall performance.
@@
Dataset Split RoboBrain (Ours) GPT-4V [2] LLaVA-OV-7B [44] InternVL2-8B [16] Qwen2-VL-7B [86] GPT-4o [68]
A12D[34] test 82.03 78.2 81.4 83.8 - 94.2
ChartQA[63] test 80.48 78.5 80 83.3 83 85.7
DocVQA[64] test 88 88.4 87.5 91.6 94.5 92.8
TextVQA[74] val 75.85 - 71.07 77.4 84.3 -
MMMU[91] val 49 56.8 48.8 51.8 54.1 69.1
MMStar[14] test 61.23 57.1 61.7 61.5 60.7 63.9
OCRBench[53] - 677 656 697 794 845 805
RealWorldQA[88] test 68.89 61.4 66.3 64.4 70.1 58.6
SeedBench[39] image 71.03 49.9 75.4 76.2 - 76.2
MMbench[55] en-dev 81.52 81.3 83.2 - - 83.4
MMbench[55] en-test 80.44 75 80.8 81.7 83 -
MME[23] test 2084 1926 1998 2210 2327 -
@@ C.1 More Results on General Benchmarks
To evaluate performance on general tasks in real-world scenarios, as is commonly done with MLLMs [2, 44, 16, 86, 68], we conducted experiments using a diverse set of image benchmarks summarized in Table 5. We leveraged the comprehensive evaluation toolkit, LMMs-Eval[95, 97], to evaluate RoboBrain’s performance on general benchmarks. These benchmarks are categorized into three classes:
  • Chart, Diagram, and Document Understanding. As key visual formats for structured OCR data, benchmarks such as AI2D [34], ChartQA [63], DocVQA [64], and OCRBench [53] were utilized. Open-source models like InternVL2-8B [16] and LLAVA-OV-7B [44] have demonstrated comparable performance to closed-source models such as GPT-4V [2]. For RoboBrain, despite being optimized primarily for multidimensional robotic tasks, it surpasses LLAVA-OV-7B [44] and GPT-4V [2] on these benchmarks, achieving a significant improvement in structured OCR tasks, with the only exceptions being DocVQA [64], where it performs slightly lower than GPT-4V [2], and OCRBench [53], where it falls slightly behind LLAVA-OV-7B [44].
  • Visual Perception and Multi-domain Reasoning. This category focuses on complex visual perception and multidisciplinary reasoning tasks. Benchmarks for visual perception include MMStar [14], MMBench [55], and MME [23], while reasoning benchmarks include MMMU [91] and SeedBench [39]. RoboBrain demonstrates comparable performance to GPT-4V [2] and LLAVA-OV-7B [44] across multiple benchmarks.
  • Real-world Understanding and Interaction. Evaluating MLLMs [2, 44, 16, 86, 68] as general-purpose assistants in real-world settings is crucial, as these scenarios extend beyond controlled environments. For this, the RealworldQA [88] benchmark was utilized. Results indicate that RoboBrain not only outperforms open-source models like LLAVA-OV-7B [44] and InternVL2-8B [16], but also exceeds closed-source models such as GPT-4V [2] and GPT-4o [68], showcasing its extensive knowledge base and strong generalization capabilities.
@@
Dataset Split / Metric RoboBrain (Ours) GPT-4V [2] LLaVA-OV-7B [44] RoboMamba [50] Qwen2-VL-7B [86]
RoboVQA[73] BLEU1 72.05 32.23 38.12 54.9 33.22
@@
OpenEQA[61] OBJECT-STATE-RECOGNITION 70.4 63.2 72.02 - 72.06
@@
To evaluate RoboBrain’s performance on robotic capabilities in real-world scenarios, we selected RoboVQA [73], OpenEQA [61], and the test set of ShareRobot, extracted from the proposed ShareRobot dataset, as robotic benchmarks for multi-dimensional assessment, as shown in Table 6. The chosen baselines include MLLMs such as GPT-4V [2], LLaVA-OV-7B [44], and Qwen2-VL-7B [86], as well as robotic models like RoboMamba [50]. Detailed descriptions of the three selected robotic benchmarks and the analysis of each results are provided below:
@@
  • OpenEQA [61] provides a robotics VQA benchmark with over 1,600 high-quality human-generated questions drawn from more than 180 real-world scenes, targeting the task of Embodied Question Answering (EQA) for environment understanding. For fairness, we evaluated all models using the prompt templates and the LLM-Score metric provided by OpenEQA [61]. Based on the evaluation results, our proposed model, RoboBrain, outperforms GPT-4V [2] overall and achieves comparable performance to other baselines. In the future, we plan to further enhance RoboBrain’s spatial intelligence to improve its generalization across scenes.
  • ShareRobot (Eval) provides a cross-scene and cross-embodiment robotics benchmark consisting of 2,050 VQA pairs, drawn from 102 diverse scenes (e.g., bedroom, laboratory, kitchen, office) and covering 12 different robot bodies. Similar to RoboVQA [73], we categorized various robotic VQA tasks into planning, planning with context, planning remaining steps, future prediction, generative affordance, past description, success (positive/negative), and discriminative affordance (positive/negative). Unlike RoboVQA benchmark [73], we utilized GPT-4o [68] to score the evaluation results instead of BLEU metrics for each task, aiming for more accurate performance assessment. Based on the results, our proposed model, RoboBrain, outperforms all baselines, demonstrating its exceptional planning capabilities across diverse scenes and embodiments.
@@ C.3 Effectiveness of ShareRobot
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 7. As shown in the table, the inclusion of ShareRobot data enhances the model’s performance compared to scenarios without ShareRobot. This highlights ShareRobot’s key role in enhancing RoboBrain’s planning capabilities.
Table 7: EExperimental results demonstrating the effectiveness of different task types. Type-1 refers to Chart, Diagram, and Document Understanding; Type-2 pertains to Visual Perception and Multi-domain Reasoning; Type-3 encompasses Real-world Understanding and Interaction. For detailed task descriptions, please refer to C.1.
@@
OneVision ShareRobot Others Type-1 Type-2 Type-3 RoboVQA[73] OpenEQA[61] ShareRobot
@@ C.4 Effectiveness of Robot Data Proportion
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 7. As shown in the table, a 4:6 ratio of robotic data is an effective choice for training, balancing performance on both the robotic and general benchmarks.
Table 8: Additional Experimental Results. “SFT Data (G:R)” indicates the ratio of training data for fine-tuning MLLMs, where “G” represents general VQA data and “R” denotes robot data (with half being ShareRobot). The total dataset size is 1.47M.
Model SFT Data(G:R) RoboVQA ShareRobot MME MMMU
LLaVA-OV-7b 6:0 36.29 27.04 2001 49.65
6:4 43.63 54.66 1945 48.83
(a) Qwen2VL-7B 6:0 24.05 28.17 2313 52.10
6:4 58.94 58.86 2295 52.33
OpenVLA-7B 6:0 4.11 21.44 1681 35.07
6:4 54.79 60.56 1722 37.25
LLaVA1.5-Qwen 6:0 24.17 26.73 1720 44.28
6:4 49.01 43.41 1732 48.33
(b) LLaVA1.5-LLaMA 6:0 21.40 25.06 1529 46.40
6:4 49.67 54.87 1722 43.41
LLaVA1.5-Vicuna 6:0 26.19 22.18 1668 30.09
6:4 50.40 51.42 1650 31.51
LLaVA1.5-Mistral 6:0 14.30 21.88 1602 23.91
6:4 36.29 57.47 1548 24.32
C.5 Different Architecture and MLLMs
To validate the effectiveness of different architecture and pre-trained MLLMs and training data in the stage 3 training setup, we selected LLaVA-OV-7B [41], OpenVLA-7B [37], and Qwen2VL-7B [86], each representing a distinct architecture among MLLMs, and conducted supervised fine-tuning (SFT) using the same proportion of training data described in the main text. As shown in Tab. LABEL:tab_additional (a), the results demonstrated that incorporating ShareRobot can significant performance improvements. For unaligned MLLMs such as LLaVA 1.5 [46] and OpenVLA, we first aligned the MLP using BLIP-558k [40] before fine-tuning. In contrast, other models that already have aligned vision encoder and LLM were directly fine-tuned.